AcmeCo reports the following in its latest quarterly report:

Authorized shares

350,000,000

Shares issued

275,500,000

Shares outstanding

262,417,205

How many shares are in the treasury stock?

If the float is 235,200,100 shares, find the number of restricted shares.

Recently, AcmeCo closed at $17.50 per share. Based on this price, find the market capitalization of the company.
Answer #1
Treasury Stock issued = No. of shares issued - No. of shares outstanding
Treasury Stock issued = 275,500,000 - 262,417,205 =13,082,795 shares
Restricted no. of shares =No. of shares outstanding - Free float no. of shares
Restricted no. of shares =262,417,205 - 235,200,100 =27,217,105
Market Cpitalization of the company =Free float no. of shares*Share price
Market Cpitalization of the company =27,217,105*$17.50 =476,299,337.50